# ElementaryNet: A Non-Strategic Neural Network for Predicting Human Behavior in Normal-Form Games

### v2

@@ 3 GameNet’s Feature Layers are Strategic
With this background, we now have a framework with which we can formally study GameNet’s feature layers.
Recall that GameNet uses its feature layers to output a predicted level-0 behavior, which then seed an iterative reasoning model in the action response layers; indeed, the feature layers are flexible enough to represent many existing level-0 models, such as the maxmax heuristic. However, this flexibility is a double-edged sword. Our first main result is that GameNet’s level-0 model can represent quantal best response to maxmax—a strategic model—to arbitrary precision.
@@
Let $q_{1}(G)=QBR_{1}(\text{maxmax}_{2}(G);1,G)$. Let $\mathcal{G}^{*}$ be the set of all games where all utilities are between $C_{max}$ and $C_{max}$, and all pairs of utilities differ by at least $C_{gap}$, where $C_{max},C_{gap}>0$ are arbitrary constants. Then, there exists an instantiation of GameNet’s feature layers that coincides with $q_{1}(G)$ for all $G\in\mathcal{G}^{*}$.
Our proof is constructive, providing parameter values for a 3-layer network that approximates this strategic model.
@@
We will first give a series of computations that are equivalent to $q_{1}(G)$ for all $G\in\mathcal{G}^{*}$, then give parameter values for GameNet’s feature layers that perform these computations. Let
@@
$E$ $=\operatorname*{relu}(U^{1}+2C_{max}B-C_{max}),\text{and}$
$Q$ $=\operatorname*{softmax}(\operatorname*{rowmax}(E)).$
First, $M_{c}$ is a matrix with each column containing the maximum utility that player 2 could realize by playing that action. $M_{*}$ is then a constant matrix containing player 2’s maxmax value. Then, assuming that all utilities differ by at least $C_{gap}$, $B$ is a matrix containing a column of ones for player 2’s maxmax action and zeros in all other columns. (Note that because all of the utilities are distinct, they have a unique maxmax action.) Assuming that all utilities are between $C_{max}$ and $C_{max}$, $E$ is a matrix containing player 1’s utilities plus $C_{max}$ in player 2’s maxmax column and zeros elsewhere. Finally, because adding a constant to player 1’s expected utilities does not affect their quantal best response, $Q$ is a vector containing player 1’s quantal best response to the maxmax action. Therefore, $Q=q_{i}(G)$.
It remains to be shown that GameNet can represent each of these computations. We will use three hidden layers, with the first two layers each containing two hidden units and the final layer containing one hidden unit; all weights and biases are zero except where specified otherwise. In the first hidden layer, we can compute $H^{1,1}=U^{1}$ by setting $w^{1}_{1,2}=1$ and $H^{1,2}=M_{c}$ by setting $w^{1}_{2,6}=1$. Then, in the second hidden layer, we can compute $H^{2,1}=U^{1}$ by setting $w^{2}_{1,1}=1$ and $H^{2,2}=B$ by setting $w^{2}_{2,2}=1/C_{gap}$, $w^{2}_{2,4}=-1/C_{gap}$, and $b^{2}_{2}=1$. Finally, in the last hidden layer, we can compute $H^{3,1}=E$ by setting $w^{3}_{1,1}=1$, $w^{3}_{1,2}=2C_{max}$, and $b^{3}_{1}=-C_{max}$. The softmax operations at the end of the feature layers complete the computation. ∎
@@ 4 A Non-Strategic Neural Network
Motivated by this failure, we next introduce a new neural network architecture and show that it is provably unable to represent strategic reasoning.
@@
In Theorem˜1, we saw that it is possible to construct non-strategic behavioral models by composing a potential function with an arbitrary response function, as long as the potential function is either dictatorial or non-encoding. Intuitively, this function composition adds an information bottleneck to the model, ensuring that some information about the agents’ utilities must be discarded before the response function can perform more complex computations.
This inspires ElementaryNet, a new neural network architecture that combines a simple, restricted potential function with a flexible feature-layer response function. ElementaryNet is a model of the form $f_{i}(G)=\sum_{p=1}^{P}w_{p}\cdot h^{p}_{i}(\Phi^{p}(G))$, where $\Phi^{p}$ are parameterized potential functions applied elementwise to the utility matrices; $h^{p}$ are parameterized response functions analogous to GameNet’s feature layers; and $w_{p}$ is a vector of probabilities. Figure˜2 illustrates the architecture.
We will study two particular instantiations of this model, differing in the specification of their potential functions. The first instantiation, which we dub the fixed-potential model, uses four linear potential functions:
@@
This choice of four potential functions is natural. Noticing that the sign of the potentials is unimportant (as the response function can easily negate all potential values before applying any other computation), these are the four distinct linear potential functions that can be constructed using the coefficients 1-1, 0, and $1$. All four also have an economic interpretation: $\varphi_{\text{own}}$ describes “single-agent” reasoning about the agent’s own payoffs; $\varphi_{\text{opp}}$ describes purely altruistic reasoning about the opponent’s payoffs; $\varphi_{\text{sum}}$ computes the welfare of each outcome; and $\varphi_{\text{diff}}$ measures the (un)fairness of each outcome. Accordingly, these potential functions can be used to represent all of the level-0 heuristics from Wright and Leyton-Brown (2019). A fixed-potential ElementaryNet model has no trainable parameters in its potential functions; its parameters consist solely of those in the response functions $h^{p}$ and the convex combination $w_{p}$.
The second instantiation, which we dub the learned-potential model, slightly loosens the restrictions on the potential functions, allowing them to be arbitrary linear functions:
$\varphi^{p}(x,y)=\theta^{p}_{x}x+\theta^{p}_{y}y,$
where the coefficients $\theta^{p}$ are trainable parameters. This model is able to represent any behavior that can be represented by a fixed-potential ElementaryNet model, but is also more flexible, being able to use other potential functions if they are justified by the training data.
@@
Our second theoretical result is that ElementaryNet is weakly non-strategic: it cannot represent quantal best-response to any dominance-responsive behavioral model.
@@
We first show that linear potential functions are either dictatorial or non-encoding. Let $\varphi(x,y)=\theta_{x}x+\theta_{y}y$ be a linear potential function. If $\theta_{x}=0$ or $\theta_{y}=0$, then $\varphi$ is dictatorial, and we are done. Otherwise, let $b>0$ be arbitrary. Let $(x_{1},y_{1})=(-b,\frac{\theta_{x}b}{\theta_{y}})$ and $(x_{2},y_{2})=(b,-\frac{\theta_{x}b}{\theta_{y}})$; we have $\varphi(x_{1},y_{1})=\varphi(x_{2},y_{2})=0$ but $|x_{1}-x_{2}|=2b>b$, as required. Similarly, let $(x_{3},y_{3})=(\frac{\theta_{x}b}{\theta_{y}},-b)$ and $(x_{4},y_{4})=(-\frac{\theta_{x}b}{\theta_{y}},b)$; we have $\varphi(x_{3},y_{3})=\varphi(x_{4},y_{4})=0$ but $|y_{3}-y_{4}|=2b>b$. Therefore, $\varphi$ is non-encoding.
Thus, ElementaryNet is a convex combinations of elementary models. The proof then follows from Theorem˜1, which shows that any such model is weakly non-strategic. ∎
Intuitively, linear functions have linear level curves, which are either axis-aligned (implying that the function is dictatorial) or extend arbitrarily far in both dimensions (implying that the function is non-encoding). In either case, linear potential functions discard a sufficient amount of information to ensure that the model cannot represent strategic reasoning. The response functions are immaterial: it is the bottleneck in the potential functions that preclude strategic behavior.
Notably, ElementaryNet does not just disagree with quantal best response on isolated games: it is incapable of representing broad categories of strategic reasoning. It can fail to be “other-responsive”, disregarding the other player’s preferences entirely; such a model cannot possibly form accurate beliefs about an opponent. Otherwise, if it is other-responsive, it can then be made to play dominated actions with probability bounded away from zero, regardless of how large the losses in utility are. We formalize these general failures to represent strategic reasoning in the appendix.
@@ 5 Experiments
While the constraints that ElementaryNet imposes on the GameNet architecture make it provably unable to represent represent strategic reasoning, only an empirical study can determine the effects this restriction in expressivity has on the model’s ability to model human behavior. Here we show the most positive result that we could hope for: that, when ElementaryNet is used as the level-0 specification for an iterative reasoning model, it nearly matches the performance of GameNet. What’s more, since our model cleanly factors descriptions of strategic and nonstrategic behavior into different parameters, we can gain insights into human behavior by varying features of the model and analyzing its parameters. We demonstrate how this type of analysis can yield interpretable results, showing that iterative reasoning is a good model of human behavior; that the precise model of iterative reasoning is relatively unimportant; that simpler level-0 specifications are inconsistent with non-strategic behavior in our data; and that there is little to gain from even richer level-0 specifications.
@@
Our dataset draws on 12 experimental studies. The first ten were used in past work performing comprehensive evaluations of behavioral game theory models (Wright and Leyton-Brown 2017, 2019); the remaining two studies, by Fudenberg and Liang (2019) and Chui, Hartline, and Wright (2023), investigated large-scale experiments on Amazon Mechanical Turk. We provide more details about these studies in the appendix.
@@
To train our models, we first randomly selected 20% of the games to use as a validation set and 20% to use as a test set, using the remaining 60% as a training set. Then, we trained up to 36 models with different hyperparameters, varying the L1 regularization coefficient applied to the neural network weights, the dropout probability, and the initial QCH model parameters. The exact hyperparameter values we used are detailed in the appendix. We report the test loss of the model that had the lowest validation loss. We used the squared L2 error between the predicted distribution and empirical distribution as our loss function, as past work has argued that it is appropriate for evaluating behavioral models (d’Eon et al. 2024). We repeated this procedure for 50 train/validation/test splits.
@@
Due to the small size of our dataset and the comparative flexibility of our models, our training and evaluation procedure had high variance, as the losses depended heavily on precisely which games were put into the training, validation, and test sets. To combat this high variance, rather than reporting losses of our models, we report the difference in loss to a reference model on the same data split. Taking paired differences in this way removes any variance caused purely by differences in the data, isolating differences in model performance. We report bias-corrected accelerated bootstrapped confidence intervals (Efron 1987) of the differences in test loss with confidence levels of both 68% and 95%.
@@
First, we compared ElementaryNet and GameNet to evaluate whether the restrictions on ElementaryNet harmed its predictive performance. We trained QCHp models using a variety of GameNet and ElementaryNet models as their level-0 input. We varied the GameNet feature layers and ElementaryNet response functions to use either 1, 2, or 3 hidden layers with widths of 10, 20, or 50. The ElementaryNet models each combined four elementary models, using each of our four fixed potential functions once. As a baseline, we also evaluated the Uniform + QCHp model (Camerer, Ho, and Chong 2004).
Our results are shown in Figure˜3. Consistent with the results reported by Hartford, Wright, and Leyton-Brown (2016), we found that the best GameNet models performed far better than the Uniform + QCHp baseline. GameNet models with only a single hidden layer performed best, while deeper models tended to overfit, achieving better training loss but worse test loss. However, we also found that the best ElementaryNet models followed closely behind. As with GameNet, ElementaryNet models having only one hidden layer performed best. In particular, the ElementaryNet model with a single hidden layer of 20 units was statistically indistinguishable from our best GameNet model at a confidence level of 5%. ElementaryNet models having a single hidden layer of either 10 or 50 units were our next-best models, each closing roughly 80% of the gap in loss between the baseline and the best GameNet model. These results show that, despite the additional constraints on ElementaryNet, it still constitutes a state-of-the-art model of human strategic behavior when used in tandem with a QCHp model.
@@
Having showed that ElementaryNet can achieve comparable performance to a state-of-the-art model, we then sought to leverage its theoretical guarantees to understand what types of reasoning might be on display in our data. We ran four follow-up experiments varying different features of the model; each one provided insight about kinds of strategic or non-strategic reasoning exhibited by the participants in our dataset.
@@
We first compared the performance of our best ElementaryNet + QCHp model to one with no QCH model at all. This experiment is similar to the one that Hartford, Wright, and Leyton-Brown (2016) ran, where they found that GameNet performed better with only a level-0 model. However, as we showed, GameNet’s level-0 model is able to emulate strategic models, making it impossible to conclude that iterative strategic reasoning is a poor model of human strategic behavior from this empirical result. In contrast, we can be sure that ElementaryNet’s level-0 model alone is non-strategic.
We trained fixed-potential ElementaryNet models with 1, 2, and 3 hidden layers, with 20 hidden units in each layer. The results, in Figure˜4, show that purely non-strategic models are an extremely poor fit to the data. The best of these models is the smallest one, with test loss comparable to that of the Uniform + QCHp baseline. The deeper models are even worse, indicating overfitting; these models are significantly worse than even the baseline model. Thus, we can conclude that iterative reasoning is indeed a good model of the behavior of our subjects, far outstripping any non-strategic model.
@@
Next, we varied the strategic model. We focused on ElementaryNet level-0 models with 1 hidden layer and the same 4 fixed potentials as before. However, we replaced the QCHp model with QCH models having arbitrary level distributions up to 1, 2, or 3 levels, learning these level distributions. The results, in Figure˜5, show that the differences between the various QCH models were relatively small. ElementaryNet models trained with QCH1 and QCH3 had performance that was statistically indistinguishable from the results with the original model using QCHp. The model using QCH2 performed the worst, but still had an average test loss far better than that of the baseline Uniform + QCHp model.
@@
Each of these QCH models have interpretable level distribution parameters, which are either discrete histograms (for the QCH1, QCH2, and QCH3 models) or Poisson distributions (for the QCHp model). Figure˜6 shows the average learned level distribution for each of the ElementaryNet models, as well as the baseline Uniform + QCHp model. These plots show a striking difference: each of the QCH models co-trained with an ElementaryNet level-0 model place over 75% of their probability on level-0 reasoners, while the baseline model places only 33% of its probability on level-0 reasoners. This difference suggests that the majority of subjects in our data are likely performing some type of relatively rich non-strategic reasoning, which the baseline model is forced to model as imperfect strategic behavior. The differences between the ElementaryNet models’ level distributions are smaller, with level-1 reasoners having approximately 20% of the probability mass for the QCH1 and QCHp models, but 13% and 10% for the QCH2 and QCH3 models, respectively.
@@
More restrictive level-0 specifications.
Having studied our strategic model extensively, we next varied features of our level-0 model. A natural question to ask is whether our nuanced definition of non-strategic behavior was necessary. After all, it would be easier to simply create a level-0 model that has access to only the agent’s own payoffs: such a model is clearly non-strategic, as it cannot reasonably form beliefs about the opponent’s behavior without knowing the opponent’s payoffs. Would this simpler level-0 specification be sufficient for modelling our subjects’ behavior?
We trained fixed-potential ElementaryNet models that only used the “own” potential, which conform to this more restrictive level-0 specification. The response functions still contained a single hidden layer with a width of 20. The results (Figure˜7, purple) show that this alternative model performed much worse than our fully-featured ElementaryNet model, with average test loss closer to that of the Uniform + QCHp baseline. This provides evidence that the experimental subjects exhibited rich non-strategic behavior, being sensitive to features like welfare or fairness that cannot be computed without knowledge of the opponent’s behavior.
Richer level-0 specifications.
Finally, we allowed our level-0 model to represent a broader set of non-strategic behavior. Rather than fixing the potential functions, we fit learned-potential ElementaryNet models, allowing the coefficients of the linear potential functions to be optimized for the training data. The results (Figure˜7, blue) show that this, too, degraded performance, causing the model to perform significantly worse than the fixed-potential benchmark. It is likely that the additional flexibility made the model more prone to overfitting, outweighing any improvements from expressing more general potential functions. This suggests that our fixed potential functions are indeed an appropriate set of basis functions for modelling non-strategic behavior.
Figure 7: ElementaryNet level-0 models with various potential functions. Overly restrictive and richer specifications of non-strategic behavior both degraded performance.

### v3

@@ 3 GameNet’s Feature Layers are Strategic
We now have a framework with which we can formally study GameNet’s feature layers. Recall that these feature layers output a predicted level-0 behavior; indeed, they are flexible enough to represent many existing level-0 models, such as the maxmax heuristic. However, this flexibility turns out to be a double-edged sword.
Our first main result is that GameNet’s level-0 model can represent a particular strategic model—quantal best response to maxmax—to arbitrary precision. We give a constructive proof, providing parameter values for a 3-layer network that approximates this strategic model.
@@
Let $q_{1}(G)=QBR_{1}(\text{maxmax}_{2}(G);1,G)$. Let $\mathcal{G}$ be the set of games where all utilities are between 0 and $C_{max}$ and all utilities differ by at least $C_{gap}$, where $C_{max},C_{gap}>0$ are arbitrary constants. Then, there exists an instantiation of GameNet’s feature layers that coincides with $q_{1}(G)$ for all $G\in\mathcal{G}$.
@@
We first give a series of computations that produce $q_{1}(G)$ for all $G\in\mathcal{G}$. Let
@@
$E$ $=\operatorname*{relu}(U^{1}+C_{max}B-C_{max}),\text{and}$
$Q$ $=\textstyle\operatorname*{softmax}([\sum_{j}E_{i,j}]_{i=1}^{n}).$
Here, $M_{c}$ is a matrix where each column contains the maximum utility that player 2 could realize by playing that action. $M_{*}$ is a constant matrix containing player 2’s maxmax value. Then, because all utilities differ by at least $C_{gap}$, $B$ is a matrix containing a column of ones for player 2’s maxmax action and zeros in all other columns. (This maxmax action is unique because all of the utilities are distinct.) Because all utilities are between 0 and $C_{max}$, $E$ is a matrix containing player 1’s utilities in player 2’s maxmax column and zeros elsewhere. Finally, $Q$ is a vector containing player 1’s quantal best response to the maxmax action. Therefore, $Q=q_{i}(G)$.
We now show that GameNet’s feature layers can represent these computations. Consider a model with three hidden layers, with two hidden units in the first two layers and one hidden unit in the final layer. Assume that all unspecified weights and biases are set to zero. In the first layer, setting $w^{1}_{1,2}=1$ gives $H^{1,1}=U^{1}$, and setting $w^{1}_{2,6}=1$ gives $H^{1,2}=M_{c}$. In the second layer, setting $w^{2}_{1,1}=1$ gives $H^{2,1}=U^{1}$, and setting $w^{2}_{2,2}=1/C_{gap}$, $w^{2}_{2,4}=-1/C_{gap}$, and $b^{2}_{2}=1$ gives $H^{2,2}=B$. In the last layer, setting $w^{3}_{1,1}=1$, $w^{3}_{1,2}=C_{max}$, and $b^{3}_{1}=-C_{max}$ gives $H^{3,1}=E$. The softmax operations at the end of the feature layers complete the proof. ∎
Note that the assumption that the games in $\mathcal{G}$ have positive utilities was made purely for clarity of exposition. It is straightforward to handle negative utilities by adding appropriate constant shifts. We prove this more general claim in Appendix B
@@ 4 A Non-Strategic Neural Network
Having seen that GameNet’s feature layers can represent strategic reasoning, we now introduce a new neural network architecture that is only capable of non-strategic behavior.
@@
In Theorem˜1, we saw that it is possible to construct non-strategic behavioral models by composing an arbitrary response function with a potential function, as long as this potential function is either dictatorial or non-encoding. Intuitively, as long as the potential function is nicely behaved, it adds an information bottleneck to the model, discarding information about the agents’ utilities before the response function can perform more complex computations.
This inspires ElementaryNet, a new architecture that adds such an information bottleneck to GameNet’s flexible feature layers. ElementaryNet is a model of the form $f_{i}(G)=\sum_{p=1}^{P}w_{p}\cdot h^{p}_{i}(\Phi^{p}(G))$, where $\Phi^{p}:\mathds{R}^{2}\to\mathds{R}$ are parameterized potential functions applied elementwise to the utility matrices; $h^{p}$ are parameterized response functions analogous to GameNet’s feature layers; and $w_{p}$ is a vector of probabilities. Figure˜2 illustrates this architecture.
We will study two particular instantiations of this model, differing in the specification of their potential functions. The first instantiation, which we dub the learned-potential model, allows the potentials to be arbitrary linear functions
$\varphi^{p}(x,y)=\theta^{p}_{x}x+\theta^{p}_{y}y,$
where the coefficients $\theta^{p}_{x}$ and $\theta^{p}_{y}$ are trainable parameters. This model is therefore able to use any linear potential function, so long as it is justified by the training data.
The second instantiation, which we dub the fixed-potential model, uses four fixed linear potential functions:
@@
This set of potential functions is natural. Noticing that the sign of the potentials is unimportant (as the response function can easily negate all potential values before applying any other computation), these are the four distinct linear potential functions that can be constructed using the coefficients 1-1, 0, and $1$. All four also have an economic interpretation. $\varphi_{\text{own}}$ describes “single-agent” reasoning about the agent’s own payoffs; $\varphi_{\text{opp}}$ describes purely altruistic reasoning about the opponent’s payoffs; $\varphi_{\text{sum}}$ computes the welfare of each outcome; and $\varphi_{\text{diff}}$ measures the (un)fairness of each outcome. Accordingly, these potential functions—specifically, $\varphi_{\text{own}}$, $\varphi_{\text{sum}}$, and $\varphi_{\text{diff}}$—can be used to represent all of the level-0 heuristics from JAIR. A fixed-potential ElementaryNet model has no trainable parameters in its potential functions; its parameters consist solely of those in the response functions $h^{p}$ and the convex combination $w_{p}$.
@@
Our second theoretical result is that ElementaryNet is weakly non-strategic: it cannot represent quantal best response to any dominance-responsive behavioral model.
@@
We will first show that any linear potential function $\varphi(x,y)=\theta_{x}x+\theta_{y}y$ is either dictatorial or non-encoding. If $\theta_{x}=0$ or $\theta_{y}=0$, then $\varphi$ is dictatorial. Otherwise, let $b>0$ be arbitrary. Let $(x_{1},y_{1})=(0,0)$ and $(x_{2},y_{2})=(2b,-2b\frac{\theta_{x}}{\theta_{y}})$. Then, we have $\varphi(x_{1},y_{1})=\varphi(x_{2},y_{2})=0$ but $|x_{1}-x_{2}|=2b>b$. Similarly, let $(x_{3},y_{3})=(2b\frac{\theta_{y}}{\theta_{x}},-2b)$; we have $\varphi(x_{1},y_{1})=\varphi(x_{3},y_{3})=0$ but $|y_{1}-y_{3}|=2b>b$. Therefore, $\varphi$ is non-encoding.
Because each of its potential functions are either dictatorial or non-encoding, ElementaryNet is therefore a convex combination of elementary models. Then, Theorem˜1 implies that it is weakly non-strategic, completing the proof. ∎
Intuitively, linear functions have linear level curves, which are either axis-aligned (implying that the function is dictatorial) or extend arbitrarily far in both dimensions (implying that the function is non-encoding). In either case, linear potential functions discard enough information to ensure that the model cannot represent strategic reasoning. The flexibility of the response functions is immaterial: it is the bottleneck in the potential functions that preclude strategic behavior.
Notably, ElementaryNet does not just disagree with quantal best response on isolated games. Instead, it is incapable of representing broad categories of strategic reasoning. It can fail to be “other-responsive”, disregarding the other player’s preferences entirely; such a model cannot possibly form accurate beliefs about an opponent. Otherwise, if it is other-responsive, it can then be made to play dominated actions with probability bounded away from zero, regardless of how large the losses in utility are. We formalize and prove these properties in Appendix C.
@@ 5 Experiments
While the constraints in ElementaryNet’s architecture make it provably unable to represent strategic reasoning, only an empirical study can determine the extent to which these constraints affect the model’s ability to model human behavior. Here we show the most positive result that we could hope for: that, when ElementaryNet is used as the level-0 specification for an iterative reasoning model, it matches the performance of GameNet. What’s more, since our model cleanly factors descriptions of strategic and nonstrategic behavior into different parameters, we can gain insights into human behavior by varying features of the model and analyzing its parameters. We demonstrate how this type of analysis can yield interpretable results, showing that iterative reasoning is a good model of human behavior; that the precise model of iterative reasoning is relatively unimportant; that far simpler level-0 specifications are inconsistent with non-strategic behavior in our data; and that level-0 specifications from prior work lose relatively little predictive power.
@@
We used a dataset consisting of results from twelve experimental studies. The first ten were used in past work comprehensively evaluating behavioral game theory models (WLB2017; JAIR). The remaining two studies ran large-scale experiments on Amazon Mechanical Turk (FL19; CHW23). In total, the dataset contains 26,553 observations across 366 distinct games. We provide more details about this data in Appendix D.
@@
To train our models, we first randomly selected 20% of the games to use as a validation set and 20% to use as a test set, using the remaining 60% as a training set. Then, we trained up to 36 models with different hyperparameters, varying the L1 regularization coefficient applied to the neural network weights, the dropout probability, and the initial QCH model parameters. The exact hyperparameter values we used are detailed in Appendix D. We report the test loss of the model that had the lowest validation loss. We used the squared L2 error between the predicted distribution and empirical distribution as our loss function, as past work has argued that it is appropriate for evaluating behavioral models (dEon2024). We repeated this procedure for 50 train/validation/test splits.
@@
Due to the small size of our dataset and the comparative flexibility of our models, the losses depend heavily on precisely which games are in the training, validation, and test sets, adding substantial variance to our training procedure. To combat this high variance, rather than reporting losses of our models, we report the difference in loss to a reference model on the same data split. Taking paired differences in this way removes any variance caused purely by differences in the data, isolating differences in model performance. We report bias-corrected accelerated bootstrapped confidence intervals (Efron1987) of the differences in test loss with confidence levels of both 68% and 95%. We also report absolute model losses in Appendix E.
@@
First, to evaluate whether ElementaryNet’s restrictions harmed its predictive performance, we compared it to GameNet. We trained QCHp models using a variety of GameNet and ElementaryNet models as their level-0 input. The ElementaryNet models each used a single learned potential function. We varied GameNet’s feature layers and ElementaryNet’s response functions to use 1, 2, or 3 hidden layers with widths of 10, 20, or 50. We also evaluated the Uniform + QCHp model as a baseline.
Our results are shown in Figure˜3. Consistent with the results reported by GameNet, we found that the best GameNet models performed far better than the Uniform + QCHp baseline. GameNet models with only a single hidden layer performed best, while deeper models had worse test loss, a likely sign of overfitting. However, the best ElementaryNet models—which also had one hidden layer—had nearly identical performance, with test losses that are statistically indistinguishable from the best GameNet model. These results show that, despite the additional constraints on ElementaryNet, it constitutes a state-of-the-art model of human strategic behavior when used in tandem with a QCHp model.
@@
Having showed that ElementaryNet + QCHp is a state-of-the-art model of strategic behavior, we then sought to leverage its theoretical guarantees to understand what types of reasoning might be on display in our data. We ran four follow-up experiments varying different features of the model. In each case, we found insights about the kinds of strategic or non-strategic reasoning exhibited by the participants in our dataset.
@@
We first compared the performance of our best ElementaryNet + QCHp model to one with no QCH model at all. This experiment is similar to the one that GameNet ran, where they found that GameNet performed better with only a level-0 model. However, we showed that GameNet’s level-0 model is able to emulate strategic models, making it impossible to conclude that iterative strategic reasoning is a poor model of human strategic behavior from this empirical result. In contrast, we can be sure that ElementaryNet’s level-0 model alone is non-strategic.
We trained learned-potential ElementaryNet models with 1, 2, and 3 hidden layers, with 50 hidden units in each layer, and no strategic model. The results (Figure˜4) show that these purely non-strategic models fit the data extremely poorly, with test losses significantly worse than even the baseline model. Thus, we can conclude that iterative reasoning is indeed a good model of the behavior of our subjects, far outstripping any non-strategic model.
@@
Next, we varied the strategic model. We focused on our best performing level-0 model: ElementaryNet with a learned potential and a response function consisting of 1 hidden layer of 50 units. However, we replaced the strategic model with QCH models having arbitrary level distributions up to 1, 2, or 3 levels, allowing these level distributions to be learned during training. The results (Figure˜5) show that the differences between the various QCH models were relatively small. ElementaryNet models trained with QCH1 and QCH3 had performance that was statistically indistinguishable from the results with the original model using QCHp. The model using QCH2 performed the worst, but still had an average test loss far better than that of the baseline Uniform + QCHp model.
@@
Each of these QCH models have interpretable level distributions, which are either discrete histograms (for the QCH1, QCH2, and QCH3 models) or Poisson distributions (for the QCHp model). Figure˜6 shows the average learned level distribution for each of the ElementaryNet models, as well as the baseline Uniform + QCHp model. These plots show that each of the QCH models co-trained with an ElementaryNet level-0 model place over 70% of their probability on level-0 reasoners, while the baseline model places only 33% of its probability on level-0 reasoners. This difference suggests that the majority of subjects in our data are likely performing some type of relatively rich non-strategic reasoning, which the baseline model is forced to model as imperfect strategic behavior. This qualitative finding echoes previous results from JAIR, who found that using level-0 models beyond the uniform baseline increased the fitted proportion of level-0 reasoners from 30% to 60%.
@@
Simpler level-0 specifications.
Having studied our strategic model extensively, we next varied features of our level-0 model. A natural question to ask here is whether our nuanced definition of non-strategic behavior was necessary. After all, it would be easier to simply create a level-0 model that makes predictions based only on the agent’s own payoffs. Such a model is clearly non-strategic, as it cannot reasonably form beliefs about the opponent’s behavior without knowing the opponent’s payoffs. Would this simpler level-0 specification be sufficient for modelling our subjects’ behavior?
We trained an ElementaryNet model with a single potential function fixed to the “own” potential, which conforms to this more restrictive level-0 specification. As before, we used a response function with one hidden layer of 50 units and a QCHp strategic model. The results (Figure˜7, light green) show that this simpler level-0 model performed far worse than our best ElementaryNet model, with average test loss closer to that of the Uniform + QCHp baseline. This provides evidence that the experimental subjects exhibited rich non-strategic behavior that cannot be expressed without knowledge of the opponent’s payoffs.
Level-0 potentials from prior work.
Lastly, we considered a middle ground: an ElementaryNet model with multiple fixed potential functions. We trained an ElementaryNet model with four potential functions, using each of the four fixed potential functions that we described in Section˜4. This model is unable to adapt its potential functions to the data, but can still express richer non-strategic behavior, such as the heuristics used by JAIR. As before, the model still used response functions with one hidden layer of 50 units and a QCHp strategic model. The results (Figure˜7, dark green) show that the model with four fixed potentials far outperformed the model with just one. However, it still did not match the performance of our best ElementaryNet model, which had just a single learned potential function. This result suggests that it may be possible to develop simple level-0 heuristic models that outperform those from past work by considering more nuanced potential functions beyond welfare and fairness.
Figure 7: ElementaryNet level-0 models with various sets of potential functions. Both alternative specifications of non-strategic behavior degraded performance.